# DISPATCH - Decentralized Informed Spatial Planning and Assignment of Tasks for Cooperative Heterogeneous Agents

### v1

@@ III Preliminaries
At inference time, agents execute their learned policies independently, using only their local observations and corresponding agent–entity graphs, without access to centralized information.
Eisenberg-Gale Model for Assignment
@@ IV EG-MARL Framework
To ensure consistency across entity types, all embeddings use a fixed-length vector equal to the landmark feature dimension. Entity-specific features that do not apply are zero-padded, allowing the GNN to process heterogeneous nodes within a unified relational graph.
@@
Workload Progress When an agent arrives and works on its assigned task, it receives a reward proportional to the workload completed during that interval. The decrease in workload follows the efficiency-driven dynamics defined in Section 2, where the rate depends on $\text{preference}_{ji}$. Accordingly, the progress reward is expressed as
@@ VI-B Evaluation Metrics
    Regret. We measure the regret of any decentralized partially observable policy $\pi$ by comparing the centralized EG objective (1a). Let $d^{*}_{ij}$ denote the shortest path distance used in the centralized EG program between agent $i$ and task $j$. Accordingly, $u_{ji}(d^{*}_{ij})$ captures utilities under the shortest centralized distances $d^{*}_{ij}$. The optimal centralized allocation is denoted by $x^{*}=(x^{*}_{ji})$, obtained from solving the centralized assignment program. The centralized objective is defined as:
@@
    Here, $u_{ji}(d^{*}_{ij})$ denotes the utility of agent $i$ serving task $j$ under the optimal centralized path $d^{*}_{ij}$, and $w_{j}$ is the task’s importance weight.
@@ VI-C Regret under Partial Observability
Following Table I, we observe that EG-MARL achieves the lowest regret for smaller teams ($N=3$), where centralized critic guidance effectively aligns local policies with the equilibrium objective. As the number of agents increases ($N=7,10$), regret rises for both frameworks, reflecting the greater difficulty of coordinating under partial observability. The Online Assignment performs competitively at larger scales, achieving lower regret than EG-MARL when $N=7$ but slightly higher when $N=10$. Overall, EG-MARL maintains more consistent proximity to the centralized optimum, while the online method demonstrates better scalability as task discovery becomes more complex.
@@ VI-D Fairness and Efficiency Analysis
Table II summarizes results across all configurations. For the Eisenberg–Gale (EG) formulation, we compare EG-MARL, Online Assignment, and the Centralized Oracle. The Hungarian and Min–Max Distance methods serve as centralized baselines without corresponding MARL or online counterparts. Metrics include completion time ($T$), total distance ($D$), and fairness ($F(\rho)$), where lower $T$ and $D$ and higher $F(\rho)$ indicate better performance.
TABLE II: Comparison across agent counts ($N=3,7,10$) for all assignment rules (Eisenberg–Gale, Hungarian, and Min–Max Distance) and algorithms (EG-MARL, Online, Centralized). Metrics include completion time $T$ and total distance $D$ (lower is better), fairness $F(\rho)$, and Jain’s Index $J(\rho)$ (higher is better).
@@
Although Table I shows that EG-MARL’s regret increases with team size, Table II reveals that it still achieves the fastest completion times for $N=7$ and $N=10$. This outcome suggests that the decentralized EG-MARL policy learns to prioritize preference alignment, completing tasks more quickly, even at the cost of longer travel distances compared to the centralized solution. The tunable parameter $\alpha$ in the distance-discounted utility $u_{ji}=(\alpha^{d_{ij}})\,\text{preference}_{ji}$ controls the balance between spatial efficiency and preference weighting. In our experiments, we set $\alpha=0.97$, favoring preference alignment and contributing to faster task completion.
@@
Table III analyzes the sensitivity of the Stochastic Online Assignment to the subset size $k$. Regret is lowest for moderate subsets (k=3k=355), which balance exploration time with the assignment optimality gap from the centralized solution. As $k$ increases, completion time $T$ decreases monotonically since larger subsets more closely approximate the centralized assignment; when $k=7$, the allocation becomes identical to the centralized EG solution, minimizing idle time but increasing travel distance. FFairness across $k$. Both fairness measures, $F(\rho)$ and Jain’s Index $J(\rho)$, remain largely stable across all subset sizes $k$, indicating that the equilibrium-based assignment preserves distributional balance regardless of how many agents are jointly considered. Minor variations at higher $k$ reflect small sensitivity but do not alter the overall fairness trend.

### v2

@@ III Preliminaries
During inference, agents execute their learned policies independently, using only their local observations, communications, and corresponding agent–entity graphs.
Eisenberg-Gale Optimization
@@ IV EG-MARL Framework
To ensure consistency across entity types, all embeddings use a fixed-length vector equal to the landmark feature dimension. Entity-specific features that do not apply are zero-padded.
@@
Workload Progress When an agent arrives and services its assigned task, it receives a reward proportional to the workload completed during that interval. The decrease in workload follows the efficiency-driven dynamics defined in Section 2, where the rate depends on $\text{preference}_{ji}$. Accordingly, the progress reward is expressed as
@@ VI-B Evaluation Metrics
    Regret. We measure the regret of any decentralized partially observable policy $\pi$ by comparing the centralized EG objective (1a). Let $d^{*}_{ij}$ denote the shortest path distance used in the centralized EG program between agent $i$ and task $j$. Accordingly, $u_{ji}(d^{*}_{ij})$ captures utilities under the shortest centralized distances $d^{*}_{ij}$. $w_{j}$ is the task’s importance weight. The optimal centralized allocation is denoted by $x^{*}=(x^{*}_{ji})$, obtained from solving the centralized assignment program. The centralized objective is defined as:
@@ VI-C Regret under Partial Observability
Following Table I, we observe that EG-MARL achieves the lowest regret for smaller teams ($N=3$). As the number of agents increases ($N=7,10$), regret rises for both frameworks, reflecting the greater difficulty of coordinating under partial observability. The Online Assignment performs competitively at larger scales, achieving lower regret than EG-MARL when $N=7$ and $N=10$. Overall, EG-MARL maintains more consistent proximity to the centralized optimum, while the online method demonstrates better scalability.
@@ VI-D Fairness and Efficiency Analysis
Table II and Figure 3 summarize results across all configurations. For the Eisenberg–Gale (EG) formulation, we compare EG-MARL, Online Assignment, and the Centralized Oracle. The Hungarian and Min–Max Distance methods serve as centralized baselines without corresponding MARL or online counterparts. Metrics include completion time ($T$), total distance ($D$), and fairness ($F(\rho)$), where lower $T$ and $D$ and higher $F(\rho)$ indicate better performance.
TABLE II: Comparison across agent counts ($N=3,7,10$) for all assignment rules and metrics include completion time $T$ and total distance $D$ (lower is better), fairness $F(\rho)$, and Jain’s Index $J(\rho)$ (higher is better).
@@
Although Table I shows that EG-MARL’s regret increases with team size, Table II reveals that it still achieves the fastest completion times for $N=7$ and $N=10$. This outcome suggests that the decentralized EG-MARL policy learns to prioritize preference alignment, completing tasks more quickly, even when this occasionally leads to slightly longer travel distances compared to the centralized solution. The tunable parameter $\alpha$ in the distance-discounted utility $u_{ji}=(\alpha^{d_{ij}})\,\text{preference}_{ji}$ controls the balance between spatial efficiency and preference weighting. In our experiments, we set $\alpha=0.97$, favoring preference alignment and contributing to faster task completion.
@@
Table III analyzes the sensitivity of the Stochastic Online Assignment to the subset size $k$.
@@
Regret is lowest for moderate subsets (k=3k=355), which balance exploration time with the assignment optimality gap from the centralized solution. As $k$ increases, completion time $T$ decreases monotonically since larger subsets more closely approximate the centralized assignment; when $k=7$, the allocation becomes identical to the centralized EG solution, minimizing idle time but increasing travel distance. In both fairness measures, $F(\rho)$ and Jain’s Index $J(\rho)$, remain largely stable across all subset sizes $k$, indicating that the online assignment preserves equilibrium balance regardless of how many agents are jointly considered. Minor variations at higher $k$ reflect small sensitivity but do not alter fairness trend.
VI-E Ablation Study
MPE Distance Penalty In the ablation study, we remove the EG fairness-shaping reward that originally guided agents toward their equilibrium-assigned tasks. Instead, the ablated reward uses a standard MPE-style distance-shaping term that is for each task $j$, the environment computes the distances from all agents to that task and subtracts the minimum distance:
$r_{\text{dist}}=-\min_{i\in\mathcal{A}}\|p_{i}-g_{j}\|.$
We also evaluate the MPE baseline reported without the workload–progress reward. In addition, we construct a second EG-MARL variant that removes the exploration reward for the ablation study. Table IV summarizes the 7-agent ablation results.
TABLE IV: Ablation results for the 7-agent setting: EG-MARL vs. Ablated (MPE reward only).
Algorithm $T$ ($\downarrow$) $D$ ($\downarrow$) $F(\rho)$ ($\uparrow$) $J(\rho)$ ($\uparrow$)
EG-MARL 2.75 10.50 1.91 0.755
Simple Spread MPE 3.62 12.08 1.895 0.737
No Exploration Reward 2.72 10.37 1.83 0.737
We observe that removing the EG fairness-shaping increases both completion time and travel distance. Despite relying solely on distance-based shaping, the ”MPE Baseline” maintains reasonable fairness, though it consistently underperforms the EG-MARL model. This indicates that the equilibrium-guided EG reward accelerates convergence while strengthening fairness.
We hypothesize that although the “No Exploration Reward” variant retains a strong navigation signal, the absence of exploration incentives leads agents to prematurely commit to early-discovered tasks. While its completion time and total distance remain comparable to those of full EG-MARL, this similarity may partly reflect statistical variability in the experimental setting rather than a true behavioral equivalence.